Circular No. 7451

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1859+226 J. Casares, P. Rodrigues, C. Zurita and T. Shahbaz, Instituto de As-

trofísica de Canarias (IAC); P. Charles and R. Hynes, Southampton University; T. Abbott, Nordic Optical Telescope (NOT); and P. Hakala, Turku Observatory, report on long-term optical photometric monitoring of the soft x-ray transient XTE J1859+226 (IAUC 7274, 7279): “R-band photometry using the IAC-80 and OGS telescopes on Tenerife and the 2.5-m NOT on La Palma showed a smooth decline of 0.017 mag/day until a large dip from R ∼ 19 to 21.5 occurred around May 30 (day 235 from the 1999 Oct. 9 outburst). The brightness recovered within five days to an extrapolation of the previous decline and then dropped again around June 11 to a minimum of R = 21.7 on June 24. Within a week it began a rapid rise into a minioutburst, reaching R ∼ 18 on July 4, since when it has been flat (with superposed variations of ∼ 0.2 mag). This is remarkably similar behavior (timescale and amplitude) to that of RX J0422+32 (Chevalier and Ilovaisky 1995, A.Ap. 297, 103). We therefore urge further coverage at all wavelengths both now and during the coming months.”

Charles, Abbott, Hakala, Casares, and also G. Israelian (IAC), report that their time-resolved (<∼ 1 min) B-band photometry with the NOT on July 5 and 6 showed large variations in the lightcurve: “On July 5 (B = 19.2, V = 18.7) the source faded by ∼ 0.3 mag for ∼ 3 hr, during which time it exhibited QPO flaring activity (up to ∼ 20 percent amplitude) on a 21.5-min period. The flaring ceased when the source returned to its predip level. The QPO timescale is very close to that reported by Hynes et al. on IAUC 7294. On July 6 no such flaring activity was seen; instead the source exhibited a smooth sinusoidal modulation (amplitude 0.1 mag) on a timescale of ∼ 6 hr, comparable to the 6.7 hr reported on IAUC 7303. This could be the orbital, or perhaps superhump modulation timescale. Further x-ray monitoring of this flaring activity is urged during the current and any future minioutbursts.”

COMETS C/1998 L2, 1998 L3, 1998 L4, 1998 L5, 1998 L6 Details on these SOHO Kreutz sungrazers are on MPEC 2000-N25.

1998 UT  α_{2000}  δ_{2000}  Comet
June 3.519  4^{h}44.^{m}5  +20^{o}36′  C/1998 L2
4.087  4 46.5  +20 40  C/1998 L3
4.936  4 49.9  +20 44  C/1998 L4
5.401  4 51.4  +20 48  C/1998 L5
5.791  4 52.9  +20 51  C/1998 L6
2000 July 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Brian G. Marsden